# A Systematic Survey of Prompt Engineering in Large Language Models: Techniques and Applications

### v1

@@ 1 Introduction
Prompt engineering has emerged as a crucial technique for enhancing the capabilities of pre-trained large language models (LLMs) and vision-language models (VLMs). It involves strategically designing task-specific instructions, referred to as prompts, to guide model output without altering parameters. The significance of prompt engineering is especially evident in its transformative impact on the adaptability of LLMs and VLMs. By offering a mechanism to fine-tune model outputs through carefully crafted instructions, prompt engineering enables these models to excel across diverse tasks and domains. This adaptability is different from traditional paradigms, where model retraining or extensive fine-tuning is often required for task-specific performance. This is the transformative promise of prompt engineering, pushing the boundaries of AI and opening doors to a future brimming with possibilities. In an ever-evolving landscape, ongoing research consistently reveals innovative approaches and applications within prompt engineering. The significance of prompt engineering is underscored by its capacity to steer model responses, enhancing the adaptability and applicability of LLMs across diverse sectors. The landscape of contemporary prompt engineering spans a spectrum of techniques, encompassing foundational methods like zero-shot and few-shot prompting to more intricate approaches such as "chain of code" prompting. The notion of prompt engineering was initially investigated and popularized in the LLMs Liu et al. (2023)Tonmoy et al. (2024)Chen et al. (2023) later extended to VLMs Wu et al. (2023)Bahng et al. (2022). Despite the extensive literature on prompt engineering within both LLMs and VLMs, a notable gap remains, particularly concerning a systematic overview of application-centric prompt engineering techniques. With recent strides in prompt engineering, there is a pressing need for a comprehensive survey that offers a nuanced understanding of applications and advancements in contemporary research. This survey dives deep into the ever-evolving landscape of prompt engineering, analyzing over 29 distinct techniques categorized by their diverse applications. Employing a systematic review approach, we meticulously delve into the intricacies of diverse cutting-edge prompting methods. Our examination encompasses their applications, the language models utilized, and the datasets subjected to experimentation, providing a detailed and nuanced analysis of the evolving landscape of prompt engineering. Additionally, we discuss the pros and cons of these techniques, offering insights into their comparative efficacy. We unveil a comprehensive taxonomy diagram that sheds light on how these techniques navigate the vast terrain of LLM capabilities. From language generation and question answering to code creation and reasoning tasks, prompt engineering empowers the LLMs into performing feats we never thought possible. By bridging the existing gap in the literature, this survey aims to serve as a valuable resource for researchers and practitioners, offering insights into the latest developments and facilitating a deeper understanding of the evolving landscape of prompt engineering. The structure of the paper is organized as follows: Section 2 presents the prompt engineering techniques from both basic to advanced by categorizing application-area and Section 3 provides a conclusion along with considerations for future research endeavors.
@@
§2.1, for tree=fill=red!50, text width=18em [ Zero-shot Prompting Radford et al. (2019), for tree=fill=pg58,text width=28em ] [ Few-shot Prompting Brown et al. (2020), for tree=fill=pg58,text width=28em ] ] [ Reasoning and Logic §2.2, for tree=fill=green!50, text width=18em [ Chain-of-Thought (CoT) Prompting Wei et al. (2022) , for tree=fill=bg4,text width=28em ] [ Automatic Chain-of-Thought (Auto-CoT) Zhang et al. (2022) , for tree=fill=bg4,text width=28em ] [ Self-Consistency Wang et al. (2022) , for tree=fill=bg4,text width=28em ] [ Logical CoT (LogiCoT) Prompting Zhao et al. (2023) , for tree=fill=bg4,text width=28em ] [ Chain-of-Symbol (CoS) Prompting Hu et al. (2023) , for tree=fill=bg4,text width=28em ] [ Tree-of-Thoughts (ToT) Prompting Yao et al. (2023a) , for tree=fill=bg4,text width=28em ] [ Graph-of-Thought (GoT) Prompting Yao et al. (2023b) , for tree=fill=bg4,text width=28em ] [ System 2 Attention Prompting Weston and Sukhbaatar (2023) , for tree=fill=bg4,text width=28em ] [ Thread of Thought (ThoT) Prompting Zhou et al. (2023) , for tree=fill=bg4,text width=28em ] [ Chain of Table Prompting Wang et al. (2024) , for tree=fill=bg4,text width=28em ] ] [ Reduce Hallucination §2.3, for tree=fill=lime!50, text width=18em [ Retrieval Augmented Generation (RAG) Lewis et al. (2020) , for tree=fill=bg32,text width=28em ] [ ReAct Prompting Yao et al. (2022) , for tree=fill=bg32,text width=28em ] [ Chain-of-Verification (CoVe) Dhuliawala et al. (2023) , for tree=fill=bg32,text width=28em ] [ Chain-of-Note (CoN) Prompting Yu et al. (2023) , for tree=fill=bg32,text width=28em ] [ Chain-of-Knowledge (CoK) Prompting Li et al. (2023d) , for tree=fill=bg32,text width=28em ] ] [ User Interaction §2.4, for tree=fill=bg25, text width=18em [ Active-Prompt Diao et al. (2023) , for tree=fill=bg8,text width=28em ] ] [ Fine-Tuning and Optimization §2.5, for tree=fill=purple!50, text width=18em [ Automatic Prompt Engineer (APE) Zhou et al. (2022) , for tree=fill=bg10,text width=28em ] ] [ Knowledge-Based Reasoning and
@@
 Paranjape et al. (2023) , for tree=fill=bg37,text width=28em ] ] [ Improving Consistency
@@
Prompting (CCoT) Chia et al. (2023) , for tree=fill=pink!50,text width=28em ] ] [ Managing Emotions and Tone §2.8, for tree=fill=bg40, text width=18em [ Emotion Prompting Li et al. (2023a) , for tree=fill=paired-light-yellow!50,text width=28em ] ] [ Code Generation and Execution §2.9, for tree=fill=cyan!50, text width=18em [ Scratchpad Prompting Nye et al. (2021) , for tree=fill=bg16,text width=28em ] [ Program of Thoughts (PoT) Prompting Chen et al. (2022),text width=28em, for tree=fill=bg16 ] [ Structured Chain-of-Thought
@@
  Li et al. (2023c),text width=28em, for tree=fill=bg16 ] [ Chain of Code (CoC) Prompting Li et al. (2023b),text width=28em, for tree=fill=bg16 ] ] [ Optimization and Efficiency §2.10, for tree=fill=bg15, text width=18em [ Optimization by Prompting Yang et al. (2023), text width=28em, for tree=fill=bg7 ] ] [ Understanding User Intent §2.11, for tree=fill=blue!50, text width=18em [ Rephrase and Respond (RaR) Prompting Deng et al. (2023), text width=28em, for tree=fill=bg38 ] ] [ Metacognition and Self-Reflection §2.12, for tree=fill=bg1, text width=18em [ Take a Step Back Prompting Zheng et al. (2023), text width=28em, for tree=fill=bg36 ] ] ]
@@ 2.1.1 Zero-Shot Prompting
Zero-shot prompting offers a paradigm shift in leveraging large LLMs. This technique Radford et al. (2019) removes the need for extensive training data, instead relying on carefully crafted prompts that guide the model toward novel tasks. Specifically, the model receives a task description in the prompt but lacks labeled data for training on specific input-output mappings. The model then leverages its pre-existing knowledge to generate predictions based on the given prompt for the new task.
@@ 2.1.2 Few-Shot Prompting
Few-shot prompting provides models with a few input-output examples to induce an understanding of a given task, unlike zero-shot prompting, where no examples are supplied Brown et al. (2020). Providing even a few high-quality examples has improved model performance on complex tasks compared to no demonstration. However, few-shot prompting requires additional tokens to include the examples, which may become prohibitive for longer text inputs. Moreover, the selection and composition of prompt examples can significantly influence model behavior, and biases like favoring frequent words may still affect few-shot results. While few-shot prompting enhances capabilities for complex tasks, especially among large pre-trained models like GPT-3, careful prompt engineering is critical to achieve optimal performance and mitigate unintended model biases.

### v2

@@ 1 Introduction
Prompt engineering has emerged as a crucial technique for enhancing the capabilities of pre-trained large language models (LLMs) and vision-language models (VLMs). It involves strategically designing task-specific instructions, referred to as prompts, to guide model output without altering parameters. The significance of prompt engineering is especially evident in its transformative impact on the adaptability of LLMs and VLMs. By offering a mechanism to fine-tune model outputs through carefully crafted instructions, prompt engineering enables these models to excel across diverse tasks and domains. This adaptability is different from traditional paradigms, where model retraining or extensive fine-tuning is often required for task-specific performance. This is the transformative promise of prompt engineering, pushing the boundaries of AI and opening doors to a future brimming with possibilities. In an ever-evolving landscape, ongoing research consistently reveals innovative approaches and applications within prompt engineering. The significance of prompt engineering is underscored by its capacity to steer model responses, enhancing the adaptability and applicability of LLMs across diverse sectors. The landscape of contemporary prompt engineering spans a spectrum of techniques, encompassing foundational methods like zero-shot and few-shot prompting to more intricate approaches such as "chain of code" prompting. The notion of prompt engineering was initially investigated and popularized in the LLMs (Liu et al., 2023)(Tonmoy et al., 2024)(Chen et al., 2023) later extended to VLMs (Wu et al., 2023)(Bahng et al., 2022). Despite the extensive literature on prompt engineering within both LLMs and VLMs, a notable gap remains, particularly concerning a systematic overview of application-centric prompt engineering techniques. With recent strides in prompt engineering, there is a pressing need for a comprehensive survey that offers a nuanced understanding of applications and advancements in contemporary research. This survey dives deep into the ever-evolving landscape of prompt engineering, analyzing over 41 distinct techniques categorized by their diverse applications. Employing a systematic review approach, we meticulously delve into the intricacies of diverse cutting-edge prompting methods. Our examination encompasses their applications, the language models utilized, and the datasets subjected to experimentation, providing a detailed and nuanced analysis of the evolving landscape of prompt engineering. Additionally, we discuss the pros and cons of these techniques, offering insights into their comparative efficacy. We present a comprehensive taxonomy diagram that illustrates how these techniques navigate the vast landscape of LLM capabilities (see Fig.2) and provide a table summarizing the datasets, employed models, and evaluation metrics (see Table1). From language generation and question answering to code creation and reasoning tasks, prompt engineering empowers the LLMs into performing feats we never thought possible. By bridging the existing gap in the literature, this survey aims to serve as a valuable resource for researchers and practitioners, offering insights into the latest developments and facilitating a deeper understanding of the evolving landscape of prompt engineering. The structure of the paper is organized as follows: Section 2 presents the prompt engineering techniques from both basic to advanced by categorizing application-area and Section 3 provides a conclusion along with considerations for future research endeavors.
@@
§2.1, for tree=fill=red!50, text width=18em [ Zero-shot Prompting (Radford et al., 2019), for tree=fill=pg58,text width=28em ] [ Few-shot Prompting (Brown et al., 2020), for tree=fill=pg58,text width=28em ] ] [ Reasoning and Logic §2.2, for tree=fill=green!50, text width=18em [ Chain-of-Thought (CoT) Prompting (Wei et al., 2022) , for tree=fill=bg4,text width=28em ] [ Automatic Chain-of-Thought (Auto-CoT) (Zhang et al., 2022) , for tree=fill=bg4,text width=28em ] [ Self-Consistency (Wang et al., 2022) , for tree=fill=bg4,text width=28em ] [ Logical CoT (LogiCoT) Prompting (Zhao et al., 2023) , for tree=fill=bg4,text width=28em ]p [ Chain-of-Symbol (CoS) Prompting (Hu et al., 2023) , for tree=fill=bg4,text width=28em ] [ Tree-of-Thoughts (ToT) Prompting (Yao et al., 2023a) , for tree=fill=bg4,text width=28em ] [ Graph-of-Thought (GoT) Prompting (Yao et al., 2023b) , for tree=fill=bg4,text width=28em ] [ System 2 Attention Prompting (Weston and Sukhbaatar, 2023) , for tree=fill=bg4,text width=28em ] [ Thread of Thought (ThoT) Prompting (Zhou et al., 2023) , for tree=fill=bg4,text width=28em ] [ Chain of Table Prompting (Wang et al., 2024) , for tree=fill=bg4,text width=28em ] [ Self-Refine Prompting (Madaan et al., 2023) , for tree=fill=bg4,text width=28em ] [ Code Prompting (Puerto et al., 2024) , for tree=fill=bg4,text width=28em ] [ Self-Harmonized CoT (ECHO) Prompting (Mekala et al., 2024) , for tree=fill=bg4,text width=28em ] [ Logic-of-Thought Prompting (Liu et al., 2024) , for tree=fill=bg4,text width=28em ] [ Instance-adaptive Prompting (IAP)  (Yuan et al., 2024) , for tree=fill=bg4,text width=28em ] [ End-to End DAG-Path (EEDP) Prompting  (Yuan et al., 2024) , for tree=fill=bg4,text width=28em ] [ Layer-of-Thoughts (LoT)  (Fungwacharakorn et al., 2024) , for tree=fill=bg4,text width=28em ] [ Narrative-of-Thought (NoT) Prompting (Zhang et al., 2024) , for tree=fill=bg4,text width=28em ] [ Buffer of Thoughts (BoT) Prompting (Yang et al., 2024) , for tree=fill=bg4,text width=28em ] [ Contrastive Denoising with Noisy Chain-of-Thought (CD-CoT)
Prompting
 (Zhou et al., 2024) , for tree=fill=bg4,text width=28em ] [ Reverse Chain-of-Thought (R-CoT) Prompting (Deng et al., 2024) , for tree=fill=bg4,text width=28em ] [ Chain of Draft (CoD) Prompting (Xu et al., 2025) , for tree=fill=bg4,text width=28em ] ] [ Reduce Hallucination §2.3, for tree=fill=lime!50, text width=18em [ Retrieval Augmented Generation (RAG) (Lewis et al., 2020) , for tree=fill=bg32,text width=28em ] [ ReAct Prompting (Yao et al., 2022) , for tree=fill=bg32,text width=28em ] [ Chain-of-Verification (CoVe) (Dhuliawala et al., 2023) , for tree=fill=bg32,text width=28em ] [ Chain-of-Note (CoN) Prompting (Yu et al., 2023) , for tree=fill=bg32,text width=28em ] [ Chain-of-Knowledge (CoK) Prompting (Li et al., 2023d) , for tree=fill=bg32,text width=28em ] ] [ User Interaction §2.4, for tree=fill=bg25, text width=18em [ Active-Prompt (Diao et al., 2023) , for tree=fill=bg8,text width=28em ] ] [ Fine-Tuning and Optimization §2.5, for tree=fill=purple!50, text width=18em [ Automatic Prompt Engineer (APE) (Zhou et al., 2022) , for tree=fill=bg10,text width=28em ] ] [ Knowledge-Based Reasoning and
@@
 (Paranjape et al., 2023) , for tree=fill=bg37,text width=28em ] ] [ Improving Consistency
@@
Prompting (CCoT) (Chia et al., 2023) , for tree=fill=pink!50,text width=28em ] ] [ Managing Emotions and Tone §2.8, for tree=fill=bg40, text width=18em [ Emotion Prompting (Li et al., 2023a) , for tree=fill=paired-light-yellow!50,text width=28em ] ] [ Code Generation and Execution §2.9, for tree=fill=cyan!50, text width=18em [ Scratchpad Prompting (Nye et al., 2021) , for tree=fill=bg16,text width=28em ] [ Program of Thoughts (PoT) Prompting (Chen et al., 2022),text width=28em, for tree=fill=bg16 ] [ Structured Chain-of-Thought
@@
  (Li et al., 2023c),text width=28em, for tree=fill=bg16 ] [ Chain of Code (CoC) Prompting (Li et al., 2023b),text width=28em, for tree=fill=bg16 ] ] [ Optimization and Efficiency §2.10, for tree=fill=bg15, text width=18em [ Optimization by Prompting (Yang et al., 2023), text width=28em, for tree=fill=bg7 ] ] [ Understanding User Intent §2.11, for tree=fill=blue!50, text width=18em [ Rephrase and Respond (RaR) Prompting (Deng et al., 2023), text width=28em, for tree=fill=bg38 ] ] [ Metacognition and Self-Reflection §2.12, for tree=fill=bg1, text width=18em [ Take a Step Back Prompting (Zheng et al., 2023), text width=28em, for tree=fill=bg36 ] ] ]
@@ 2.1.1 Zero-Shot Prompting
Zero-shot prompting offers a paradigm shift in leveraging large LLMs. This technique removes the need for extensive training data, instead relying on carefully crafted prompts that guide the model toward novel tasks (Radford et al., 2019). Specifically, the model receives a task description in the prompt but lacks labeled data for training on specific input-output mappings. The model then leverages its pre-existing knowledge to generate predictions based on the given prompt for the new task.
@@ 2.1.2 Few-Shot Prompting
Few-shot prompting provides models with a few input-output examples to induce an understanding of a given task, unlike zero-shot prompting, where no examples are supplied (Brown et al., 2020). Providing even a few high-quality examples has improved model performance on complex tasks compared to no demonstration. However, few-shot prompting requires additional tokens to include the examples, which may become prohibitive for longer text inputs. Moreover, the selection and composition of prompt examples can significantly influence model behavior, and biases like favoring frequent words may still affect few-shot results. While few-shot prompting enhances capabilities for complex tasks, especially among large pre-trained models like GPT-3, careful prompt engineering is critical to achieve optimal performance and mitigate unintended model biases.
@@ 2.2.10 Chain-of-Table Prompting
2.2.11 Self-Refine Prompting
Self-Refine prompting, proposed by  Madaan et al. (2023), enhances LLM performance by iteratively refining outputs through self-generated feedback, mimicking human revision. While LLMs can handle a wide range of tasks, they often struggle with complex objectives, ambiguous goals, or multi-step reasoning, leading to initial responses with inaccuracies or flawed logic. Inspired by human iterative refinement, Self-Refine enables LLMs to improve their outputs through a structured three-step process: generating an initial response, prompting the model to critique its own output, and refining the response based on this feedback. This cycle continues until predefined stopping criteria are met, allowing the model to produce more accurate and contextually relevant results. Unlike traditional prompting methods, which rely solely on a single-step response, Self-Refine fosters incremental improvement, making it particularly effective for tasks requiring nuanced reasoning. Experimental results demonstrate significant performance gains, with GPT-4 improving by 8.7 points in code optimization, 13.9 points in code readability, and 21.6 points in sentiment reversal tasks, showcasing its potential to enhance the reasoning and adaptability of LLMs across various domains.
2.2.12 Code Prompting
Pre-training on code enhances the reasoning capabilities of LLMs, yet the underlying mechanisms driving this improvement remain poorly understood. To investigate this, Puerto et al. (2024) examines the impact of input representation on LLM reasoning, specifically exploring whether reformulating natural language (NL) problems into code can trigger conditional reasoning abilities. This led to the introduction of Code Prompting, a technique that reformulates NL tasks into structured code, enabling direct prompting of text+code LLMs without relying on external code execution. Experiments on three reasoning benchmarks, ConditionalQA, BoardgameQA, and ShARC, demonstrate that code prompts significantly outperform traditional text-based prompts. On average, GPT 3.5 achieved a performance gain of 8.42 F1 score, while Mistral showed an average improvement of 4.22 across the three datasets.
2.2.13 Self-Harmonized Chain-of-Thought (ECHO) Prompting
While Chain-of-Thought prompting enhances reasoning in LLMs, methods like Auto-CoT, which automate demonstration generation, face challenges from misleading similarity (incorrect rationales in similar examples) and ineffective diversity (irrelevant or overly varied patterns). To address these issues, Mekala et al. (2024) introduced ECHO, a self-harmonized prompting framework that unifies diverse reasoning paths into a coherent pattern, balancing automation with robustness. ECHO operates through three key stages: (1) Question Clustering, where Sentence-BERT embeddings and k-means group questions into clusters; (2) Demonstration Sampling, which selects representative questions from each cluster and generates rationales using Zero-Shot-CoT; and (3) Demonstration Unification, where rationales are iteratively refined using a dynamic prompting mechanism to align reasoning patterns. This process minimizes diversity-induced noise while retaining adaptability. ECHO surpassed Auto-CoT by an average of 2.8% across 10 reasoning benchmarks (arithmetic, commonsense, symbolic) while demonstrating greater efficiency. It retained performance with 50% fewer examples, showing only a -0.8% dip compared to Few-Shot-CoT’s -1.3% decline. The method also achieved 2.3% gains over Auto-CoT in Mixtral-8x7B, though it remained behind GPT-3.5, a gap attributed to differences in the quality of reasoning rationales.
2.2.14 Logic-of-thought Prompting
LLMs often exhibit unfaithful reasoning, where the generated conclusions diverge from the intermediate reasoning steps. Logic-of-Thought prompting (Liu et al., 2024) is a neuro-symbolic framework developed to mitigate this issue by enriching prompts with logical information derived from propositional logic. LoT operates in three phases: (1) Logic Extraction, during which LLMs identify propositions and logical relationships from input texts; (2) Logic Extension, in which a Python-based module applies formal logical laws (e.g., contraposition) to infer additional expressions; and (3) Logic Translation, where the extended logic is rendered back into natural language and appended to the original prompt to ensure contextual fidelity. Moreover, Logic-of-thought is designed to integrate seamlessly with other prompting strategies such as CoT, Self-Consistency, and ToT prompting. Reported evaluations indicate that Logic-of-thought can improve CoT accuracy on the ReClor benchmark by 4.35%, enhance CoT prompting with Self-Consistency on LogiQA by 5%, and further boost ToT prompting performance on the ProofWriter dataset by 8%. Additionally, by preserving natural language representations throughout the process, Logic-of-Thought avoids the symbolic extraction errors that can impair other neuro-symbolic systems, such as SatLM.
2.2.15 Instance-adaptive Prompting (IAP)
Yuan et al. (2024) tackle the generalization constraints of static task-level prompts (e.g., "Let’s think step by step") in zero-shot CoT reasoning by introducing Instance-Adaptive Prompting (IAP), a saliency-driven framework designed to dynamically tailor prompts to individual instances. Through information flow analysis of attention layers, the authors identified distinct patterns: effective reasoning correlates with strong semantic flow from questions to prompts in shallow layers and from integrated question-prompt representations to rationales in deeper layers. In contrast, fragmented or weak flows are indicative of suboptimal reasoning performance. IAP optimizes reasoning fidelity through two adaptive strategies. The first, IAP-ss (Sequential Substitution), enhances efficiency by iteratively testing prompts until predefined saliency thresholds are met. The second, IAP-mv (Majority Vote), prioritizes robustness by aggregating saliency scores across multiple prompts to determine consensus answers. Empirical evaluations underscore the broad applicability of IAP: in mathematical reasoning tasks (GSM8K, SVAMP), IAP-mv boosts the performance of LLaMA-3-8B and Qwen-14B by +1.82% and +3.31%, respectively, compared to static prompts. It achieves 19.25% accuracy on causal judgment tasks, outperforming baselines at 16.04%, and surpasses Self-Discover by +21.7% on MMLU commonsense reasoning with Qwen-14B.
2.2.16 End-to End DAG-Path (EEDP) Prompting
End-to-End DAG-Path (EEDP) prompting (Hong et al., 2024) addresses the limitations of traditional graph-flattening methods, such as adjacency lists and edge lists, which struggle with long-distance reasoning in graph-related tasks for LLMs. EEDP’s key insight is that conventional flattened representations often lose critical long-range dependencies essential for effective reasoning. To mitigate this, EEDP prioritizes the main backbone paths connecting graph endpoints (nodes with zero in-degree or out-degree) while preserving adjacency lists to maintain local contextual information. The EEDP framework operates through three key stages: (1) preprocessing input graphs into directed acyclic graphs (DAGs) using breadth-first search (BFS) to eliminate cycles, (2) extracting hierarchical paths between endpoints, and (3) compressing shared path segments with a differential pointer algorithm, effectively reducing token length by 55% on molecular graphs. EEDP was evaluated on tasks such as Edge Prediction Connectivity Prediction (EPCP) and Edge Prediction Distance Prediction (EPDP) using educational (Merged_1000) and molecular (ZINC_test_2500) datasets. The evaluation results highlighted significant performance gains over traditional baselines, with EPCP showing a +10.21% accuracy improvement on Merged_1000 and +16.76% on ZINC_test_2500. Similarly, EPDP achieved a +4.73% accuracy boost on Merged_1000 and an impressive +30.13% on ZINC_test_2500.
2.2.17 Layer-of-Thoughts (LoT) Prompting
LLMs demonstrate strong performance in many reasoning tasks yet frequently face challenges with the precision–recall trade-off and explainability, particularly in complex legal retrieval scenarios. Layer-of-Thoughts (LoT) prompting (Fungwacharakorn et al., 2024) introduces a hierarchical framework that leverages constraint hierarchies to structure the reasoning process, thereby enhancing both retrieval accuracy and interpretability. In the context of legal document retrieval, LoT organizes reasoning into "layer thoughts" (conceptual stages) and "option thoughts" (partial solutions), applying sequential constraints to iteratively filter and refine candidate responses. For instance, the framework employs a three-layer process: (1) a Keyword Filtering Layer (KFL) that extracts LLM-generated keywords to initially filter documents using metrics such as at-least-k; (2) a Semantic Filtering Layer (SFL) that prioritizes documents based on multi-level relevance criteria and aggregation metrics; and (3) a Final Confirmation Layer (FCL) that validates the remaining candidates against the original query. By integrating both hard constraints (required) and soft constraints (preferential), LoT not only delivers explainable reasoning but also outperforms state-of-the-art models, for example, achieving an F2 score of 0.835 (with precision of 0.838 and recall of 0.839) on Japanese Civil Law retrieval compared to 0.807 for JNLP, and reaching near-perfect recall (0.966) in German traffic law contexts.
2.2.18 Narrative-of-Thought (NoT) Prompting
Temporal reasoning remains a significant challenge for LLMs, particularly in inferring global temporal relationships from unordered events. To evaluate this capability, Zhang et al. (2024) introduced Temporal Graph Generation (TGG), a benchmark designed to assess LLMs’ proficiency in constructing directed acyclic graphs (DAGs) representing event timelines. Experimental results revealed that smaller LLMs (<10B) lagged behind GPT-3.5/4 by approximately 50%, with even GPT-4 facing difficulties due to alignment constraints. To overcome these limitations, the authors proposed Narrative-of-Thought (NOT), a prompting strategy that enhances temporal reasoning without requiring additional model training. NOT comprises three core components: (1) Structural Representation, where events are encapsulated in a Python class and processed through code completion; (2) NOT Prompting template, which generates temporally grounded narratives to guide the construction of temporal graphs; and (3) Narrative-Aware Demonstrations, utilizing GPT-4-generated few-shot examples optimized for both conciseness and accuracy. Results demonstrated that NOT significantly improves the performance of small LLMs, with LLaMA3-8B achieving an F1 score of 42.2, closely matching GPT-3.5’s 45.7, while exhibiting superior structural coherence.
2.2.19 Buffer of Thoughts (BoT) Prompting
Existing prompting methods often struggle to balance universality, efficiency, and robustness in complex reasoning. To address this, Yang et al. (2024) introduced Buffer of Thoughts (BoT), a framework that enhances LLMs through reusable high-level reasoning patterns. BoT overcomes the limitations of single-query methods (e.g., manual exemplar reliance) and multi-query approaches (e.g., computational inefficiency) by introducing a meta-buffer that distills "thought-templates" from diverse tasks and a dynamic buffer-manager that continuously refines them as new problems are solved. BoT retrieves task-specific thought-templates (e.g., structured problem-solving approaches) and adaptively instantiates them, mimicking human analogical reasoning to eliminate manual prompt design and recursive exploration. Experiments across 10 benchmarks demonstrate its state-of-the-art performance, achieving gains of 11% on Game of 24, 20% on Geometric Shapes, and 51% on Checkmate-in-One, while using just 12% of the computational cost of multi-query methods like Tree-of-Thoughts. Notably, BoT enhances smaller models, with Llama3-8B + BoT surpassing Llama3-70B in accuracy, showing its potential to democratize efficient reasoning at scale.
2.2.20 Contrastive Denoising with Noisy Chain-of-Thought (CD-CoT) Prompting
Contrastive Denoising with Noisy Chain-of-Thought (CD-CoT) (Zhou et al., 2024) addresses the challenge of "noisy rationales" in chain-of-thought prompting, where irrelevant or incorrect intermediate reasoning steps degrade LLM performance. The NoRa (Noisy Rationales) dataset highlights this issue, showing that LLMs often perform worse with flawed rationales than with no examples at all, as they tend to mimic incorrect reasoning. Existing methods like self-correction and self-consistency offer limited solutions, as self-correction fails without external feedback, and self-consistency selects frequent answers without resolving reasoning flaws. CD-CoT mitigates this by contrasting noisy rationales with clean ones, rephrasing flawed examples, selecting optimal reasoning paths, and voting on the most consistent answer. Experiments show that CD-CoT improves accuracy by 17.8% on average, significantly outperforming baselines and enhancing LLMs’ robustness in reasoning-intensive tasks.
2.2.21 Reverse Chain-of-Thought (R-CoT) Prompting
Deng et al. (2024) introduced the Reverse Chain-of-Thought (R-CoT) pipeline, a novel approach to enhancing geometric reasoning in LMMs by addressing dataset limitations such as low quality, diversity, and fidelity. R-CoT operates in two stages: GeoChain, which generates high-fidelity geometric images with detailed step-by-step descriptions of geometric relationships (e.g., midlines, radii), and Reverse A&Q, which derives questions from reasoning chains using LLMs, ensuring accurate multi-step problem generation. By prioritizing answer-aware question synthesis, R-CoT mitigates visual hallucinations and reasoning errors in LMMs. The resulting GeoMM dataset includes 20 geometric shapes categorized by complexity, incorporating relational questions often missing in existing datasets like MAVIS and GeomVerse. GeoMM combines high-fidelity images with diverse Q&A pairs, enriched by geometric theorems and line operations. Experimental results demonstrate that R-CoT-trained models achieve state-of-the-art performance, with the 8B-parameter model surpassing GPT-4o by 12.5% on MathVista and 14.5% on GeoQA, while smaller models (2B, 7B) also set new benchmarks.
2.2.22 Chain of Draft (CoD) Prompting
Chain of Draft (CoD) (Xu et al., 2025), a novel prompting strategy designed to enhance efficiency in complex reasoning tasks. Unlike traditional CoT prompting, which emphasizes detailed step-by-step reasoning, CoD generates concise, information-dense outputs at each step, mirroring human problem-solving strategies where only essential insights are noted. While CoT improves reasoning accuracy, it often leads to verbose outputs and increased computational costs. CoD mitigates this by constraining word usage in each reasoning step, reducing latency and token consumption without sacrificing accuracy. This efficiency-oriented approach is particularly valuable for real-world applications where computational resources and response time are critical. Experimental results across arithmetic, commonsense, and symbolic reasoning benchmarks show that CoD matches or even outperforms CoT in accuracy while significantly lowering token usage and latency. In some cases, CoD achieved comparable accuracy with an 80% reduction in output tokens as well as an average latency reduction of 76.2%, demonstrating its potential as a lightweight yet effective alternative to traditional prompting strategies.